karen is 6 years older than nina five years from now the sum of their ages will be 52 how old are both of them
Answer by ankor@dixie-net.com(22740)

karen is 6 years older than nina five years from now the sum of their ages will be 52 how old are both of them
:
let k = Karen's age
let n = Nina's
:
Write an equation for each statement
:
" karen is 6 years older than nina"
k = n + 6
" five years from now the sum of their ages will be 52"
(k + 5) + (n + 5) = 52
k + n + 10 = 52
k + n = 52 - 10
k + n = 42
Replace k with (n+6), from the first statement
(n + 6) + n = 42
2n = 42 - 6
2n = 36
n = 36/2
n = 18 is Nina's age
then
k = 18 + 6
k = 24 is Karen's age
:
:
Check this in the statement
" five years from now the sum of their ages will be 52"
24+5 + 18+5 =
29 + 23 = 52